OpenStudy (anonymous):

how do you solve this question? When x>1, what expression has a value less then x?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you have some choices? i mean \[x-1\] for example is always less than x

OpenStudy (anonymous):

and if \[x>1\] then 0 is less than x. as is 1.

OpenStudy (anonymous):

a. 2.5xsquared b. xcubed+4 c.squareroot of x-2 d.squareroot of x+x

OpenStudy (anonymous):

how would you solve this?

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

well what i would do is just check numbers greater than 1 and see what works

OpenStudy (anonymous):

says x must be greater than 1, so lets check the first one with some number bigger than 1

OpenStudy (anonymous):

\[2.5x^2\] if x is 10 gives \[2.5\times 100=250\] which is bigger than 10 so no

OpenStudy (anonymous):

\[x^3+4\] is not worth checking because in your head you see if you plug in say 10 you get a much bigger number back

OpenStudy (anonymous):

\[\sqrt{x-2}\] will work although x has to be bigger than 2, not 1. say x = 11 then \[\sqrt{11-2}=\sqrt{9}=3\] which is less than 11 so pick that one